OpenStudy (anonymous):

A bag contains four chips, numbered 3, 5, 7, and 8. Two chips are drawn at random. What is the probability that the sum of the two chips is an odd number?

OpenStudy (anonymous):

I'm not sure about this but it seems logic.. the options that can be drawn are: 3+5=8 3+7=10 3+8=11 5+7=12 5+8=13 7+8=15 3 of these are odd and 3 are even so 50%.. However... I'm not completely sure if this shouldn't be solved with the change of the 8 chip being picked

OpenStudy (anonymous):

Thanks~